# A Continuous-Time Ensemble Kalman–Bucy Smoother for Causal Inference and Model Discovery

### 28 Apr 2026 v1

@@ 3 EnKBS for State Estimation, Causal Inference, and Model Identification
In all experiments, we use the Euler–Maruyama scheme for numerical time integration [69]. For the backward smoothing pass, we discretize the formulation (9) and propagate the ensemble backward in time.
@@ 3.1 State Estimation in a Relatively High-Dimensional Chaotic System: The Lorenz-96 Model
The first example considers the state estimation of the Lorenz-96 system. Introduced by Edward Lorenz, it was originally designed to mimic the propagation of atmospheric waves on a midlatitude zonal circle [18, 70]. It is now widely used as a benchmark test case in data assimilation. Let $\mathbf{x}=(x_{1},\dots,x_{n})^{\top}\in\mathbb{R}^{n}$. We consider the stochastic Lorenz-96 dynamics:
@@
The quadratic term $(x_{j+1}-x_{j-2})x_{j-1}$ mimics nonlinear advection and preserves the total energy, which is proportional to $\sum_{i}x_{i}^{2}$. The linear term $x_{j}$ represents mechanical or thermal dissipation and provides system damping, and the constant forcing $F$ drives the system into a chaotic regime when it exceeds a certain threshold value. With this parameter choice, the deterministic system has $13$ positive and one neutral Lyapunov exponent, corresponding to the dimension of the unstable-neutral subspace [63, 70].
@@ 3.1.1 The Experiment Setup
In this test, we observe every other state component (the even indices) and treat the remaining components as hidden [55]. In other words, the hidden signal process (1a) and observational process (1b) correspond respectively to
@@
We set the observation rate to be equal to the numerical integration time step $\Delta t=0.005$. A reference trajectory $\mathbf{x}_{\mathrm{ref}}$ is generated over the time window $[0,T]$ with $T=100$, and a segment is shown in Figure 1. We initialize the ensemble by adding small perturbations to the initial reference state. For localization, we use the periodic distance between indices $i$ and $j$:
@@ 3.1.2 Results
$\mathrm{RMSE}:=\sqrt{\frac{1}{nK}\sum_{k=1}^{K}\left\|\overline{\mathbf{x}}(t_{k})-\mathbf{x}_{\mathrm{ref}}(t_{k})\right\|^{2}},$
where $K$ is the total number of time steps and $\overline{\mathbf{x}}$ denotes the ensemble mean (filter or smoother). The results show that, with localization and mild inflation, the smoother consistently improves upon the filter on Lorenz-96. While a suitable inflation factor slightly reduces the error, localization has a pronounced impact on the stability and convergence of EnKBS. Compared with EnKBF, EnKBS is more sensitive to ensemble size. Without localization, EnKBS typically requires $m\gtrsim 25$ to remain stable in this setting; for smaller $m$, the backward pass diverges. In contrast, EnKBF remains stable even with ensembles as small as $m=3$. With localization, EnKBS remains stable for much smaller ensembles (down to $m=5$) and still outperforms the corresponding filter estimate (not shown). Figure 2 compares the time trajectories of $x_{1}$ between the reference solution and the EnKBF/EnKBS mean, with parameters $r_{0}=3$ and $\delta^{2}=1.005$.
$\delta^{2}\backslash r_{0}$ 1 2 3 4 5 8 15 18
1 0.714 0.662 0.656 0.660 0.668 0.704 0.734 1.025
1.0001 0.712 0.662 0.656 0.659 0.668 0.704 0.733 1.025
1.001 0.700 0.659 0.654 0.658 0.666 0.703 0.732 1.026
1.005 0.683 0.664 0.654* 0.654 0.661 0.698 0.728 1.033
1.01 NaN 0.762 0.683 0.667 0.671 0.698 0.727 1.041
1.02 NaN NaN NaN NaN NaN 0.848 1.037 1.062
@@ 3.2 Causal Inference in a Nonlinear Dyad Model
We incorporate the EnKBS into the recently developed assimilative causal inference (ACI) framework [16]. Unlike traditional approaches based solely on time-averaged statistics or model forecasts, ACI casts causal inference as an inverse problem through Bayesian data assimilation, enabling time-dependent diagnosis of causal dynamics. In this case study, we use a nonlinear dyad model to demonstrate that the EnKBS-ACI framework is able to infer the causal intensity and temporal influence range of a hidden driver from observations of its effect, dynamically identifying the instantaneous cause-and-effect role interplay.
The key idea behind ACI is that, if a variable $\mathbf{y}$ drives another $\mathbf{x}$, then conditioning on the effect $\mathbf{x}$’s future evolution should retrospectively improve the estimate of the driver $\mathbf{y}$’s current state. In ACI, the instantaneous causal influence intensity is quantified by an information-gain metric (ACI metric) based on the relative entropy between the smoother distribution and the real-time filtering distribution. By further evaluating the information gain across lagged smoothers of different assimilating time windows, we can also obtain the (forward) causal influence range (CIR) of a relationship, which characterizes how far the causal signal propagates in time. This CIR is objective, i.e., it does not rely on subjective or empirical threshold choices. Here, using the EnKBS to approximate the filter and smoother distributions removes the need to directly observe the driver for causal inference, as long as its effect is observed.
@@ 3.2.1 A Nonlinear Dyad Model with Extreme Events
Here, $W_{u}$ and $W_{v}$ are independent standard Wiener processes. This physics-constrained low-order model produces extreme events through a trigger–feedback cycle: $v$ modulates the damping of $u$, while $u$ feeds back onto $v$ [71]. Specifically, the growth rate of $u$ is controlled by the time-varying coefficient $d_{u}+cv$. When intermittent growth of $v$ exceeds the threshold $d_{u}/c$, the system enters an anti-damping regime and $u$ grows rapidly. As $u$ amplifies, the term $cu^{2}$ suppresses $v$, which restores damping in $u$ and terminates the burst in its amplitude, ultimately conserving the energy of the system.
3.3 The Experiment Setup
The following parameter values are adopted in (17):
@@ 3.3 The Experiment Setup
We consider both causal directions, the triggering direction $v\to u$ and the feedback direction $u\to v$. For $v\to u$, the filter and smoother estimate the hidden state $v$ given the observations of $u$. Notice that when $u$ is observed, the $v$ dynamics (17b) become conditionally linear and Gaussian, fitting the conditional Gaussian nonlinear system (CGNS) structure [72, 73]. This provides an analytic filter and smoother that enables a benchmark comparison with the EnKBF and EnKBS [49, 16]. For $u\to v$, we estimate the hidden state $u$ from observations of $v$. Note that this predictability arises through the coupling term $cu^{2}$ in (17b). However, because the map $u\mapsto u^{2}$ is non-injective, conditioning on $v$ alone does not resolve the sign of $u$. As a result, when the true $u$ is close to 0, ensemble members may split into opposite branches, artificially inflating the variance and increasing the estimation error. Hence, we fix a reference solution of length $T=30$ over which $u$ does not change sign for both directions. See the top panel of Figure 3. We observe that inflation does not noticeably change the results in this regime, and we therefore do not apply inflation.
@@ 3.3.1 Results
Triggering: $v\rightarrow u$}
Figure 4 compares the RMSE between the EnKBF/EnKBS and CGNS analytic filter/smoother estimates of $v$ as a function of ensemble size $m$. As $m$ increases, the EnKBS ensemble mean and covariance converge to the analytic statistics, with the resulting ACI diagnostics converging accordingly. Figures 3 and 5 show the ACI results with an ensemble size of $m=50$ and $m=10$ respectively. With $m=50$, the EnKBS diagnostics closely match the analytic benchmark. We therefore display only the CIR and ACI metric results with the EnKBS. For $m=10$, the ACI metric does not preserve relative peak magnitudes and underestimates the CIR at the beginning of the rise of $v$, i.e., during the build-up phase of $u$’s extreme event; this is especially pronounced at prolonged or delayed onsets. Nevertheless, it captures the peak locations and yields qualitatively similar patterns for the CIR and ACI metric.
@@
Feedback: $u\rightarrow v$}
In this regime, experiments use an ensemble size of $m=50$ (based on the results of Figure 4). Figure 6 shows the estimation errors and empirical standard deviations for both the filter and the smoother. During extreme events, the smoother provides a clear correction relative to the filter. No sign-splitting is observed since $u>0$ throughout.
@@
To interpret these patterns, given a time series of $\mathrm{d}v$ in the $v$-equation (17b), the filter infers $u$ through the negative coupling term $cu^{2}$. As a result, the filter cannot anticipate the growth and peak of $u$ unless the subsequent decline of $v$ is observed. In contrast, the smoother can trace back from this decline to correctly estimate $u$. Consequently, both the CIR and the ACI metric become large during the onset stage, and the CIR gradually shortens as the decline of $v$ is imminent. Later, $v$ starts to decay. This phenomenon already provides an informative signal to the filter, and the information flow $u\rightarrow v$ has largely been realized and manifested in the CIR and ACI metric during the onset period. Therefore, the filter can achieve an accurate estimate via dynamical interpolation, so the additional future information becomes negligible. As a result, the ACI metric does not persist after the peak and through the full demise of the extreme events.
@@ 3.4 Data-Driven Discovery of Dynamics with EnKBS-Based Sampling
In this subsection, we adapt a causality-based learning algorithm to recover the underlying dynamics of partially observed systems. The key idea is to formulate model learning as an iterative procedure that alternates among conditional sampling, structure identification, and parameter estimation [15]. The procedure consists of three steps:
@@
We incorporate physics constraints, e.g., energy conservation in quadratic terms, to prevent nonphysical finite-time blow-up and stabilize learning [74, 75]. These constraints are enforced in the parameter-estimation step together with other constraints to facilitate learning convergence.
@@ 3.4.1 The Lorenz-84 System
Numerical tests are performed on the Lorenz-84 model [76]:
@@ 3.4.2 The Experiment Setup
This library matches the CGNS structure, enabling direct comparison with the CGNS optimal sampling benchmark [15, 72]. In the structure-identification step, the constant term is always included by default.
We use the following initial model guess, same as [15]:
@@ 3.4.3 Results
To quantify structure convergence, we introduce a causation entropy indicator matrix $\mathbf{C}\in\{0,1\}^{N\times M}$ with system dimension $N=3$ and the number of candidate functions $M$ [15]. Each entry is obtained by thresholding the causation entropy:
@@
where $\mathrm{CE}_{i,j}$ denotes the causation entropy from the $j$-th candidate function to the $i$-th state equation [77, 78]. We then compute the Frobenius norm of the matrix difference $\|\mathbf{C}-\mathbf{C}_{\mathrm{stable}}\|_{F}$, where $\mathbf{C}_{\mathrm{stable}}$ denotes the indicator matrix after the identified structure has stabilized to the true model structure (18). The decay in Figure 8(b) indicates that the model structure is fully identified after about 10 iterations.
@@
$\theta^{x}_{x}$ $\theta^{y}_{y}$ $\theta^{z}_{z}$ $\theta^{x}_{y^{2}}$ $\theta^{x}_{z^{2}}$ $\theta^{y}_{xz}$
Truth -0.25 -1 -1 -1 -1 -4
Identified -0.2459 -1.0125 -1.0037 -1.0023 -1.0011 -3.9732
$\theta^{y}_{xy}$ $\theta^{z}_{xy}$ $\theta^{z}_{xz}$ $\theta^{x}_{1}$ $\theta^{y}_{1}$ $\theta^{z}_{1}$
Truth 1 4 1 2 1 0
Identified 1.0023 3.9732 1.0011 1.9946 1.0088 -0.0061
Table 3: Coefficients in the true Lorenz-84 system (18) and the identified model at the 120th iteration. The identified model has the same structure as the truth. Here, $\theta^{x}_{xy}$ denotes the coefficient of the feature xyxy in the $x$-equation, and analogous notation is used for other terms.
@@ 4 Conclusions and Discussion
Through numerical experiments, we show that the EnKBS consistently improves the filtered estimates by reducing error and uncertainty, converging asymptotically to the optimal smoother when an analytic benchmark is available. We further demonstrate that covariance localization and inflation are essential for numerical stability in high-dimensional settings. With regularization, spurious long-range correlations and rank deficiency are mitigated, preventing numerical divergence in high dimensions. Beyond state estimation, we integrate EnKBS into a Bayesian causal inference framework and a causality-based iterative learning algorithm for partially observed systems [16, 15]. With an ensemble size of $O(10)$, EnKBS provides effective state estimates and trajectory samples of unobserved variables in nonlinear systems without analytic smoothers, enabling causality diagnostics around extreme events and supporting model discovery.
As a final remark, although measurements are typically available at discrete times, the continuous-time EnKBS provides a complementary perspective to discrete-time and mixed-time smoothers [53, 57]. This viewpoint is consistent with the continuous-time nature of the underlying physics, and is particularly natural when the goal is to discover dynamical mechanisms. We interpret observed variables as components of the same continuous-time system, rather than as external analysis updates. Assimilation then exploits intrinsic coupling and information transfer within the system, using observed trajectories to constrain the inferred hidden states.
Future work is twofold: expanding the EnKBS methodology and its dynamical discovery applications. On the methodology side, while EnKBS performs well on test systems in this paper, these examples largely involve quadratic or conditionally quadratic structures. A natural next step is to evaluate EnKBS on more realistic, high-dimensional, and strongly nonlinear models. In particular, we can study accuracy and convergence as well as filter/smoother degeneracy together with possible mitigation strategies. On the application side, further experiments and analyses are needed to clarify the mechanisms behind causal information flow and causal reversals in driver-feedback cycles. For example, alongside the forward CIR, we can investigate with the EnKBS the backward CIR to assess the causal precursors that influenced the currently observed event, enabling attribution of significant events in general complex systems [31].
Acknowledgment
The research of N.C. is funded by the Army Research Office W911NF-23-1-0118. Z.J and M.A. are partially supported as research assistants under this grant. M.A. is also supported in part by NSF Award DMS-2023239. The work of S.R. has been partially funded by Deutsche Forschungsgemeinschaft (DFG) – Project-ID 318763901 – SFB1294.

### 03 May 2026 v2

@@ 2 EnKBS for State Estimation, Causal Inference, and Model Identification
In all experiments, we use the Euler–Maruyama scheme for numerical time integration (Kloeden and Platen, 1992). For the backward smoothing pass, we discretize the formulation Eq. (9) and propagate the ensemble backward in time.
@@ 2.1 State Estimation in a Relatively High-Dimensional Chaotic System: The Lorenz-96 Model
The first example considers the state estimation of the Lorenz-96 system. Introduced by Edward Lorenz, it was originally designed to mimic the propagation of atmospheric waves on a midlatitude zonal circle (Lorenz, 2006; Lorenz and Emanuel, 1998). It is now widely used as a benchmark test case in data assimilation. Let $\vec{x}=(x_{1},\dots,x_{n})^{\top}\in\mathbb{R}^{n}$. We consider the stochastic Lorenz-96 dynamics:
@@
The quadratic term $(x_{j+1}-x_{j-2})x_{j-1}$ mimics nonlinear advection and preserves the total energy, which is proportional to $\sum_{i}x_{i}^{2}$. The linear term $x_{j}$ represents mechanical or thermal dissipation and provides system damping, and the constant forcing $F$ drives the system into a chaotic regime when it exceeds a certain threshold value. With this parameter choice, the deterministic system has $13$ positive and one neutral Lyapunov exponent, corresponding to the dimension of the unstable-neutral subspace (Bocquet and Carrassi, 2017; Lorenz and Emanuel, 1998).
@@ 2.1.1 The Experiment Setup
In this test, we observe every other state component (the even indices) and treat the remaining components as hidden (Yang and Delsole, 2009). In other words, the hidden signal process Eq. (1a) and observational process Eq. (1b) correspond respectively to
@@
We set the observation rate to be equal to the numerical integration time step $\Delta t=0.005$. A reference trajectory $\vec{x}_{\mathrm{ref}}$ is generated over the time window $[0,T]$ with $T=100$, and a segment is shown in Figure 1. We initialize the ensemble by adding small perturbations to the initial reference state. For localization, we use the periodic distance between indices $i$ and $j$:
@@ 2.1.2 Results
$\mathrm{RMSE}:=\sqrt{\frac{1}{nK}\sum_{k=1}^{K}\left\|\overline{\vec{x}}(t_{k})-\vec{x}_{\mathrm{ref}}(t_{k})\right\|^{2}},$
where $K$ is the total number of time steps and $\overline{\vec{x}}$ denotes the ensemble mean (filter or smoother). The results show that, with localization and mild inflation, the smoother consistently improves upon the filter on Lorenz-96. While a suitable inflation factor slightly reduces the error, localization has a pronounced impact on the stability and convergence of EnKBS. Compared with EnKBF, EnKBS is more sensitive to ensemble size. Without localization, EnKBS typically requires $m\gtrsim 25$ to remain stable in this setting; for smaller $m$, the backward pass diverges. In contrast, EnKBF remains stable even with ensembles as small as $m=3$. With localization, EnKBS remains stable for much smaller ensembles (down to $m=5$) and still outperforms the corresponding filter estimate (not shown). Figure 2 compares the time trajectories of $x_{1}$ between the reference solution and the EnKBF/EnKBS mean, with parameters $r_{0}=3$ and $\delta^{2}=1.005$.
@@ 2.2 Causal Inference in a Nonlinear Dyad Model
We incorporate the EnKBS into the recently developed assimilative causal inference (ACI) framework (Andreou et al., 2026). Unlike traditional approaches based solely on time-averaged statistics or model forecasts, ACI casts causal inference as an inverse problem through Bayesian data assimilation, enabling time-dependent diagnosis of causal dynamics. In this case study, we use a nonlinear dyad model to demonstrate that the EnKBS-ACI framework is able to infer the causal intensity and temporal influence range of a hidden driver from observations of its effect, dynamically identifying the instantaneous cause-and-effect role interplay.
The key idea behind ACI is that, if a variable $\vec{y}$ drives another $\vec{x}$, then conditioning on the effect $\vec{x}$’s future evolution should retrospectively improve the estimate of the driver $\vec{y}$’s current state. In ACI, the instantaneous causal influence intensity is quantified by an information-gain metric (ACI metric) based on the relative entropy between the smoother distribution and the real-time filtering distribution. By further evaluating the information gain across lagged smoothers of different assimilating time windows, we can also obtain the (forward) causal influence range (CIR) of a relationship, which characterizes how far the causal signal propagates in time. This CIR is objective, i.e., it does not rely on subjective or empirical threshold choices. Here, using the EnKBS to approximate the filter and smoother distributions removes the need to directly observe the driver for causal inference, as long as its effect is observed.
@@ 2.2.1 A Nonlinear Dyad Model with Extreme Events
Here, $W_{u}$ and $W_{v}$ are independent standard Wiener processes. This physics-constrained low-order model produces extreme events through a trigger–feedback cycle: $v$ modulates the damping of $u$, while $u$ feeds back onto $v$ (Chen and Majda, 2020). Specifically, the growth rate of $u$ is controlled by the time-varying coefficient $d_{u}+cv$. When intermittent growth of $v$ exceeds the threshold $d_{u}/c$, the system enters an anti-damping regime and $u$ grows rapidly. As $u$ amplifies, the term $cu^{2}$ suppresses $v$, which restores damping in $u$ and terminates the burst in its amplitude, ultimately conserving the energy of the system.
2.2.2 The Experiment Setup
The following parameter values are adopted in Eq. (17):
@@ 2.2.2 The Experiment Setup
We consider both causal directions, the triggering direction $v\to u$ and the feedback direction $u\to v$. For $v\to u$, the filter and smoother estimate the hidden state $v$ given the observations of $u$. Notice that when $u$ is observed, the $v$ dynamics Eq. (17b) become conditionally linear and Gaussian, fitting the conditional Gaussian nonlinear system (CGNS) structure (Chen and Majda, 2018; Liptser and Shiryayev, 1978). This provides an analytic filter and smoother that enables a benchmark comparison with the EnKBF and EnKBS (Andreou and Chen, 2024a; Andreou et al., 2026). For $u\to v$, we estimate the hidden state $u$ from observations of $v$. Note that this predictability arises through the coupling term $cu^{2}$ in Eq. (17b). However, because the map $u\mapsto u^{2}$ is non-injective, conditioning on $v$ alone does not resolve the sign of $u$. As a result, when the true $u$ is close to 0, ensemble members may split into opposite branches, artificially inflating the variance and increasing the estimation error. Hence, we fix a reference solution of length $T=30$ over which $u$ does not change sign for both directions. See the top panel of Figure 3. We observe that inflation does not noticeably change the results in this regime, and we therefore do not apply inflation.
@@ 2.2.3 Results
Triggering: $v\rightarrow u$}. Figure 4 compares the RMSE between the EnKBF/EnKBS and CGNS analytic filter/smoother estimates of $v$ as a function of ensemble size $m$. As $m$ increases, the EnKBS ensemble mean and covariance converge to the analytic statistics, with the resulting ACI diagnostics converging accordingly. Figures 3 and 5 show the ACI results with an ensemble size of $m=50$ and $m=10$ respectively. With $m=50$, the EnKBS diagnostics closely match the analytic benchmark. We therefore display only the CIR and ACI metric results with the EnKBS. For $m=10$, the ACI metric does not preserve relative peak magnitudes and underestimates the CIR at the beginning of the rise of $v$, i.e., during the build-up phase of $u$’s extreme event; this is especially pronounced at prolonged or delayed onsets. Nevertheless, it captures the peak locations and yields qualitatively similar patterns for the CIR and ACI metric.
@@
Feedback: $u\rightarrow v$}. In this regime, experiments use an ensemble size of $m=50$ (based on the results of Figure 4). Figure 6 shows the estimation errors and empirical standard deviations for both the filter and the smoother. During extreme events, the smoother provides a clear correction relative to the filter. No sign-splitting is observed since $u>0$ throughout.
@@
To interpret these patterns, given a time series of $\mathrm{d}v$ in the $v$-equation Eq. (17b), the filter infers $u$ through the negative coupling term $cu^{2}$. As a result, the filter cannot anticipate the growth and peak of $u$ unless the subsequent decline of $v$ is observed. In contrast, the smoother can trace back from this decline to correctly estimate $u$. Consequently, both the CIR and the ACI metric become large during the onset stage, and the CIR gradually shortens as the decline of $v$ is imminent. Later, $v$ starts to decay. This phenomenon already provides an informative signal to the filter, and the information flow $u\rightarrow v$ has largely been realized and manifested in the CIR and ACI metric during the onset period. Therefore, the filter can achieve an accurate estimate via dynamical interpolation, so the additional future information becomes negligible. As a result, the ACI metric does not persist after the peak and through the full demise of the extreme events.
@@ 2.3 Data-Driven Discovery of Dynamics with EnKBS-Based Sampling
In this subsection, we adapt a causality-based learning algorithm to recover the underlying dynamics of partially observed systems. The key idea is to formulate model learning as an iterative procedure that alternates among conditional sampling, structure identification, and parameter estimation (Chen and Zhang, 2023). The procedure consists of three steps:
@@
We incorporate physics constraints, e.g., energy conservation in quadratic terms, to prevent nonphysical finite-time blow-up and stabilize learning (Majda and Harlim, 2012b; Harlim et al., 2014). These constraints are enforced in the parameter-estimation step together with other constraints to facilitate learning convergence.
@@ 2.3.1 The Lorenz-84 System
Numerical tests are performed on the Lorenz-84 model (Lorenz, 1984):
@@ 2.3.2 The Experiment Setup
This library matches the CGNS structure, enabling direct comparison with the CGNS optimal sampling benchmark (Chen and Zhang, 2023; Chen and Majda, 2018). In the structure-identification step, the constant term is always included by default.
We use the following initial model guess, same as (Chen and Zhang, 2023):
@@ 2.3.3 Results
To quantify structure convergence, we introduce a causation entropy indicator matrix $\mathbf{C}\in\{0,1\}^{N\times M}$ with system dimension $N=3$ and the number of candidate functions $M$ (Chen and Zhang, 2023). Each entry is obtained by thresholding the causation entropy:
@@
where $\mathrm{CE}_{i,j}$ denotes the causation entropy from the $j$-th candidate function to the $i$-th state equation (Elinger and Rogers, 2021; Sun and Bollt, 2014). We then compute the Frobenius norm of the matrix difference $\|\mathbf{C}-\mathbf{C}_{\mathrm{stable}}\|_{F}$, where $\mathbf{C}_{\mathrm{stable}}$ denotes the indicator matrix after the identified structure has stabilized to the true model structure Eq. (18). The decay in Figure 8(b) indicates that the model structure is fully identified after about 10 iterations.
@@
Table 3: Coefficients in the true Lorenz-84 system Eq. (18) and the identified model at the 120th iteration. The identified model has the same structure as the truth. Here, $\theta^{x}_{xy}$ denotes the coefficient of the feature xyxy in the $x$-equation, and analogous notation is used for other terms.
[…]
@@
Through numerical experiments, we show that the EnKBS consistently improves the filtered estimates by reducing error and uncertainty, converging asymptotically to the optimal smoother when an analytic benchmark is available. We further demonstrate that covariance localization and inflation are essential for numerical stability in high-dimensional settings. With regularization, spurious long-range correlations and rank deficiency are mitigated, preventing numerical divergence in high dimensions. Beyond state estimation, we integrate EnKBS into a Bayesian causal inference framework and a causality-based iterative learning algorithm for partially observed systems (Andreou et al., 2026; Chen and Zhang, 2023). With an ensemble size of $O(10)$, EnKBS provides effective state estimates and trajectory samples of unobserved variables in nonlinear systems without analytic smoothers, enabling causality diagnostics around extreme events and supporting model discovery.
As a final remark, although measurements are typically available at discrete times, the continuous-time EnKBS provides a complementary perspective to discrete-time and mixed-time smoothers (Kalman and Bucy, 1961; Bergemann and Reich, 2012). This viewpoint is consistent with the continuous-time nature of the underlying physics, and is particularly natural when the goal is to discover dynamical mechanisms. We interpret observed variables as components of the same continuous-time system, rather than as external analysis updates. Assimilation then exploits intrinsic coupling and information transfer within the system, using observed trajectories to constrain the inferred hidden states.
Future work is twofold: expanding the EnKBS methodology and its dynamical discovery applications. On the methodology side, while EnKBS performs well on test systems in this paper, these examples largely involve quadratic or conditionally quadratic structures. A natural next step is to evaluate EnKBS on more realistic, high-dimensional, and strongly nonlinear models. In particular, we can study accuracy and convergence as well as filter/smoother degeneracy together with possible mitigation strategies. On the application side, further experiments and analyses are needed to clarify the mechanisms behind causal information flow and causal reversals in driver-feedback cycles. For example, alongside the forward CIR, we can investigate with the EnKBS the backward CIR to assess the causal precursors that influenced the currently observed event, enabling attribution of significant events in general complex systems (Andreou and Chen, 2025).
[…]
All MATLAB code used to generate the numerical results and figures reported in this paper is available at https://github.com/jiangzh67/EnKBS under the MIT License.